OpenStudy (anonymous):

Suppose a function is defined as f(x) = [(x + a)(x + b)] divided by [(x + c)(x + d)]. What are its zeros, asymptotes, and holes? Now write a rational function with numbers instead of a, b, c, and d. Give the function's zeros, asymptotes, and holes. (Note that more than one rational function can have the same zeros, asymptotes, and holes. Therefore, many answers to this question are possible.)

terenzreignz (terenzreignz):

Okay... zeros of a function (in particular, a rational one, such as this) would be the values of x that would make the NUMERATOR equal to zero. the numerator here is (x+a)(x+b) If we set (x+a)(x+b) = 0 What are the values of x?

OpenStudy (anonymous):

1?

terenzreignz (terenzreignz):

No....

OpenStudy (anonymous):

-x?

terenzreignz (terenzreignz):

Okay... simply put (x+a)(x+b) = 0 only means either x+a = 0 or x+b = 0 Solve both these equations...

OpenStudy (anonymous):

oh okay x+a-a=x x+b-b=x Like that?

terenzreignz (terenzreignz):

=0 at the right sides, and not = x

OpenStudy (anonymous):

Oh okay

terenzreignz (terenzreignz):

Okay, like this... \[\large x+a = 0\] subtract a from both sides \[\large x+a\color{red}{-a}= 0 \color{red}{-a}\] simplify \[\large x = -a\] tadaa\ :D

OpenStudy (anonymous):

Thanks! :D

terenzreignz (terenzreignz):

That's not all of the answers yet...

terenzreignz (terenzreignz):

Just one of them.

OpenStudy (anonymous):

Oh okay, I'm listening. Go on

terenzreignz (terenzreignz):

I really have to go now... I'm sure there are others here that can help you... The other equation is x+b=0 Would you please solve this?

terenzreignz (terenzreignz):

(Not forgetting of course, that as of now, we are only solving for the zeros of the function)

OpenStudy (anonymous):

x+b-b=0

terenzreignz (terenzreignz):

you forgot to do the same to the right-side

OpenStudy (anonymous):

0+b-b=0?

terenzreignz (terenzreignz):

uhh... no... check the way I solved the x+a = 0 again

OpenStudy (anonymous):

OH! x+b-b=0-b

terenzreignz (terenzreignz):

Yes... continue...

OpenStudy (anonymous):

x=-b

terenzreignz (terenzreignz):

That's good. So, x = -b and x = -a are the zeros of the function. To find the asymptotes, this time, set the denominator (x+c)(x+d) to be equal to zero. Solve (x+c)(x+d) = 0

OpenStudy (anonymous):

x+c-c=0-c x=-c x+d-d=0-d x=-d

terenzreignz (terenzreignz):

That's good :) So, your asymptotes are x = -c and x = -d Next, the holes... They are the values which would make BOTH the numerator and denominator equal to zero. The zeros of the numerator are -a and -b The zeros of the denominator are -c and -d Do they have any in common?

OpenStudy (anonymous):

they are negative?

terenzreignz (terenzreignz):

No, I mean, does the numerator have any zeros that are also zeros of the denominator?

OpenStudy (anonymous):

is it x?

terenzreignz (terenzreignz):

No... what are the zeros of the numerator again?

OpenStudy (anonymous):

-a and -b

terenzreignz (terenzreignz):

And are any of these two also zeros of the denominator?

OpenStudy (anonymous):

no

terenzreignz (terenzreignz):

Then the function has no holes :P

OpenStudy (anonymous):

ohh okay